# From Text to Emoji: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs

### v2

@@ Abstract
As the demand for human-like interactions with LLMs continues to grow, so does the interest in manipulating their personality traits, which has emerged as a key area of research. Methods like prompt-based In-Context Knowledge Editing (IKE) and gradient-based Model Editor Networks (MEND) have been explored but show irregularity and variability. IKE depends on the prompt, leading to variability and sensitivity, while MEND yields inconsistent and gibberish outputs. To address this, we employed Opinion QA Based Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), to manipulate the Big Five personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. After PEFT, models such as Mistral-7B-Instruct and Llama-2-7B-chat began generating emojis, despite their absence in the PEFT data. For instance, Llama-2-7B-chat generated emojis in 99.5% of extraversion-related test instances, while Mistral-7B-Instruct did so in 92.5% of openness-related test instances. Explainability analysis indicated that the LLMs used emojis intentionally to express these traits. This paper provides a number of novel contributions. First, introducing an Opinion QA dataset for PEFT-driven personality manipulation; second, developing metric models to benchmark LLM personality traits; third, demonstrating PEFT’s superiority over IKE in personality manipulation; and finally, analysing and validating emoji usage through explainability methods such as mechanistic interpretability and in-context learning explainability methods.
1 Introduction and Related Work
@@
As Large Language Models (LLMs) become more integral across various industries, there is increasing interest in enhancing not only their linguistic capabilities but also their ability to exhibit personality traits (Hilliard et al. (2024) Hu et al. (2024) Serapio-García et al. (2023) Dan et al. (2024)). Psychological research has shown that personality traits significantly influence human communication, including tone and verbosity (Liu and Sun (2020) Kennison et al. (2024)), raising the question of whether LLMs can be manipulated to exhibit similar expressions to make their communication more nuanced and adaptable. Several frameworks, such as the Dark Triad (Jonason and Webster (2010)), the 16 Personality Factors (Cattell and Mead (2008)), and the Big Five Traits (Gosling et al. (2003)), have been used to analyse LLM personality. Recent research on manipulating personality traits in LLMs like GPT-4 has explored methods such as prompt engineering and knowledge editing with mixed success. While GPT-3 and similar models can exhibit traits through prompts, results are often inconsistent due to prompt dependency (Miotto et al. (2022); Jiang et al. (2023b); Li et al. (2022); Caron and Srivastava (2022)). Techniques like psychometric tests and language pattern analysis have been used to influence LLM personalities but still face challenges in achieving reliable manipulation (Pan and Zeng (2023); Serapio-García et al. (2023); La Cava et al. (2024). Li et al. (2023)) introduced Unsupervisedly-Built Personalized Lexicons (UBPL) to adjust Big Five traits during decoding, avoiding fine-tuning but risking training data bias. Similarly, Weng et al. (2024) and Dan et al. (2024) proposed ControlLM and P-tailor to efficiently simulate traits using control vectors, though these methods add complexity and may struggle with scalability. Finally, Mao et al. (2023) employed knowledge editing techniques like IKE, MEND, SERAC, and Prompt to manipulate traits like agreeableness, but MEND and SERAC still yield inconsistent results.
This paper addresses the challenges of personality manipulation in LLMs by introducing a novel approach grounded in the Big Five personality model. We present a new Opinion QA dataset and methodologies for systematically adjusting personality traits in LLMs. Utilising Quantized Low-Rank Adaptation (QLoRA), a method within Parameter-Efficient Fine-Tuning (PEFT) (Dettmers et al. (2023a)), we demonstrate that LLMs can achieve more consistent and enduring personality expressions. This approach enhances model adaptability in interactions and reveals new behaviours, such as spontaneous emoji generation in Mistral-7B-Instruct and Llama-7B-chat, absent in the original models, following PEFT-based adjustments. Through mechanistic interpretability (Bereska and Gavves (2024)), we confirmed that pre-training data likely influenced emoji usage, while ICL (Brown (2020)) explainability verified that the emojis were not random artifacts. Our findings suggest this phenomenon represents a novel mode of expression linked to specific personality traits (Figure 1), introducing a new dimension of LLM communication that integrates verbal and visual elements which enhances user engagement, improves emotional expressiveness in digital assistants, and enables more personalized user experiences in areas such as mental health, education, and customer service (Votintseva et al. (2024)).
@@ 2 Methodology
This paper explores manipulation in autoregressive transformer models using Parameter-Efficient Fine-Tuning (PEFT) and In-Context Knowledge Editing (IKE), focusing on Llama-2-7B-Chat (Touvron et al. (2023)), Llama-3-8B-Instruct (MetaAI (2023)), and Mistral-7B-Instruct (Jiang et al. (2023a)).
Personality Dataset and Metric Classifier
@@
This work expands on the dataset from Mao et al. (2023), which consisted of opinion-based QA on specific topics, by better aligning with the topics and more accurately capturing personality traits. This study adds Openness and Conscientiousness to the original traits of Extraversion, Agreeableness, and Neuroticism. While Mao et al. (2023) excluded these dimensions, considering them similar to Agreeableness in generating viewpoints, we argue their inclusion is vital for a comprehensive analysis and understanding of trait influence on opinions. The dataset contains $5000$ instances, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait).  [^1] Dataset: https://huggingface.co/datasets/holistic-ai/personality_manipulation. A GPT-3.5-based model generated opinion texts using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression. The generated text was analysed using word clouds and text analysis to identify key linguistic patterns, thematic elements, and ensure lexical diversity associated with the Big Five personality traits, as detailed in Appendix A.1. The text was manually verified to ensure alignment with intended traits. A multi-class personality classifier based on RoBERTa (Liu et al. (2019)), fine-tuned on the enhanced dataset, achieved $91.9\%$ accuracy on the test set.  [^2] Classifier: https://huggingface.co/holistic-ai/personality_classifier. Classifier validation involved human verification of textual feature importance using SHAP (Lundberg and Lee (2017)) and LIME (Ribeiro et al. (2016)) to ensure predictions aligned with human understanding. For instance, in the sentence, "I believe the First Indochina War had its consequences, paving the way for the withdrawal of French colonial forces. However, there were many factors at play, and it’s important to acknowledge the contributions of everyone involved." certain words play a pivotal role in predicting the Agreeableness trait as observed in Figures 2 and 3. Terms such as "contributions", "acknowledge", and "believe" have a strong positive contribution to the prediction of Agreeableness, as they suggest inclusiveness, recognition, and a conciliatory tone, which are characteristic of Agreeableness. On the other hand, words like "consequences" contribute negatively to the Agreeableness prediction because it often connotes conflict, repercussions, or negative outcomes, whereas Agreeableness is characterized by cooperation, empathy, and a focus on harmony and positive social interactions (Liu and Sun (2020)).
@@
Thus, as seen above and in other examples inA.3, the results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity. Through above analysis, we can conclude that the personality classifier has successfully captured the expected patterns from perspective of feature attribution. Further details about the classifier are in A.2.

### v3

@@ Abstract
The manipulation of the personality traits of large language models (LLMs) has emerged as a key area of research. Methods like prompt-based In-Context Knowledge Editing (IKE) and gradient-based Model Editor Networks (MEND) have been explored but show irregularity and variability; IKE depends on the prompt, leading to variability and sensitivity, while MEND yields inconsistent and gibberish outputs. To address this, we employed Opinion QA Based Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), to manipulate the Big Five personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. After PEFT, models such as Mistral-7B-Instruct and LLaMA-2-7B-chat began generating emojis, even though no emojis were present in the PEFT data. For instance, LLaMA-2-7B-chat generated emojis in 99.5% of extraversion-related test instances, while Mistral-7B-Instruct did so in 92.5% of openness-related test instances. ICL Explainability analysis indicated that the LLMs used emojis intentionally to express these traits. Mechanistic Interpretability analysis showed that this latent behaviour of LLMs could be traced to specific neurons that became activated or amplified after PEFT. This paper provides a number of novel contributions. First, introducing an Opinion QA dataset for PEFT-driven personality manipulation; second, developing metric models to benchmark LLM personality traits; third, demonstrating PEFT’s superiority over IKE in personality manipulation; and finally, analysing and validating emoji usage through explainability methods such as Mechanistic Interpretability and In-context learning Explainability methods.
From Text to Emoji: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs
Navya Jain1,2, Zekun Wu1,2thanks: Corresponding author: zekun.wu@holisticai.com, Cristian Munoz1, Airlie Hilliard1, Xin Guan1 Philip Treleaven2, Emre Kazim1, Adriano Koshiyama1 1Holistic AI, 2University College London
1 Introduction
@@
As Large Language Models (LLMs) become more integral across various industries, there is increasing interest in enhancing not only their capabilities like reasoning, planning, comprehension, but also their ability to exhibit personality traits Hilliard et al. (2024); Hu et al. (2024); Serapio-García et al. (2023); Dan et al. (2024). Psychological research has shown that personality traits significantly influence human communication, including tone and verbosity Liu and Sun (2020); Kennison et al. (2024), raising the question of whether LLMs can be manipulated to exhibit similar expressions to make their communication more nuanced and adaptable. Several frameworks, such as the Dark Triad Jonason and Webster (2010), the 16 Personality Factors Cattell and Mead (2008), and the Big Five Personality Model Gosling et al. (2003), have been used to analyse LLMs’ personality. Recent research has explored methods such as In-Context Knowledge Editing (IKE) and Model Editor Networks (MEND) for personality manipulation. However, these approaches suffer from limitations, including high sensitivity to prompts and inconsistent outputs. This paper addresses the challenges of personality manipulation in LLMs by introducing a novel approach grounded in the Big Five personality model. We present a new Opinion QA dataset and methodologies for systematically adjusting personality traits in LLMs. Utilising Quantized Low-Rank Adaptation (QLoRA), a method within Parameter-Efficient Fine-Tuning (PEFT) Dettmers et al. (2023a), we demonstrate that LLMs can achieve more consistent and enduring personality expressions.
@@
This approach enhances model adaptability in interactions and reveals new behaviours, such as spontaneous emoji generation in Mistral-7B-Instruct and LLaMA-7B-Chat, absent in the original models, following PEFT-based manipulations. ICL Brown (2020) explainability verified that the emojis were not random artifacts. Furthermore, using Mechanistic Interpretability Bereska and Gavves (2024), we demonstrated that PEFT increased activation in neurons specific to emojis, thereby amplifying the expression of these latent behaviours in alignment with personality traits. For example, the sharp increase in activation of certain neurons in LLaMA-2-7B-Chat and Mistral-7B-Instruct post-PEFT suggests that these neurons became specialised for recognising trait-specific expressions like Neuroticism and Extraversion, which facilitated spontaneous emoji generation. Our findings suggest this phenomenon represents a novel mode of expression linked to specific personality traits (Figure 1), introducing a new dimension of LLM communication that integrates verbal and visual elements which enhances user engagement, improves emotional expressiveness in digital assistants, and enables more personalized user experiences in areas such as mental health, education, and customer service Votintseva et al. (2024).
2 Related Work
This research is grounded in three key areas: LLM fine-tuning, mechanistic interpretability, and personality manipulation. In this section, we review the related work in these domains, focusing on fine-tuning techniques, mechanistic interpretability, and personality manipulations in LLMs.
2.1 Parameter-Efficient Fine-Tuning: LoRA and QLoRA
Low-Rank Adaptation (LoRA) Hu et al. (2021) is a Parameter-Efficient Fine-Tuning method that reduces memory requirements by freezing the original pre-trained model weights and introducing trainable low-rank matrices into the model. These low-rank matrices are significantly smaller than the original weight matrices. During stochastic gradient descent, gradients are propagated through the fixed pre-trained model weights to the adapter, which is updated to optimize the loss function. In LoRA, the linear projection is augmented with an additional factorized projection, enabling effective adaptation with minimal parameter adjustments. Mathematically, Given a projection
$\mathbf{Y}=\mathbf{XW}+s\mathbf{XL}_{1}\mathbf{L}_{2}$ (1)
where $\mathbf{X}\in\mathbb{R}^{b\times h}$, $\mathbf{W}\in\mathbb{R}^{h\times o}$, $\mathbf{L}_{1}\in\mathbb{R}^{h\times r}$, $\mathbf{L}_{2}\in\mathbb{R}^{r\times o}$, and $s$ is a scaling factor. Building on LoRA’s efficiency, Quantized Low-Rank Adaptation (QLoRA) Dettmers et al. (2024) has recently been introduced to further enhance the scalability of fine-tuning large models by reducing memory usage without sacrificing performance. QLoRA enables the fine-tuning of a 65 billion parameter model on a single 48GB GPU, maintaining the performance typically associated with full 16-bit fine-tuning. This is achieved through 4-bit NormalFloat (NF4) quantization, which utilizes Quantile Quantization Dettmers et al. (2021) to ensure an even distribution of values across quantization bins, and Double Quantization, which reduces memory usage by compressing the quantization constants. Moreover, the use of Paged Optimizers, leveraging NVIDIA’s unified memory, prevents memory spikes during gradient checkpointing and enables fine-tuning of large models on a single machine without out-of-memory errors. Thus, QLoRA not only retains the benefits of LoRA’s parameter-efficient adaptation but also introduces quantization strategies that make it feasible to fine-tune large-scale models with significantly lower hardware requirements.
2.2 Mechanistic Interpretability and Neuron Analysis
Mechanistic interpretability Olah et al. (2020) involves analyzing a model’s fundamental components—such as features, neurons, layers, and connections—to understand its internal operational mechanics. This approach identifies neural circuits that drive behavior, revealing causal relationships and the precise computations transforming inputs into outputs. In this context, features are considered as the fundamental units of representation, with each feature being a human-interpretable property encoded in model activations Rai et al. (2024). Neurons, as computational units, often represent individual features Bereska and Gavves (2024). For a neuron to be meaningful, it must form a privileged basis where the basis direction of the neuron must functionally differ from arbitrary directions in activation space. Non-linear activation functions privilege the basis directions defined by neurons, making individual neuron analysis a meaningful approach Elhage et al. (2022) and a practical method for uncovering network functionality Dai et al. (2021); Voita et al. (2023). Neuron analysis can be further categorized into five major methods: visualization, corpus-based, probing-based, causation-based, and miscellaneous techniques, each providing unique insights into network functionality and contributing to a holistic understanding of the model’s internal operations Sajjad et al. (2022).
2.3 Personality Manipulation
Recent research on manipulating personality traits in LLMs like GPT-4 has explored methods such as prompt engineering and knowledge editing with mixed success. While GPT-3 and similar models can exhibit traits through prompts, results are often inconsistent due to prompt dependency Miotto et al. (2022); Jiang et al. (2023b); Li et al. (2022); Caron and Srivastava (2022). Techniques like psychometric tests and language pattern analysis have been used to influence LLM personalities but still face challenges in achieving reliable manipulation Pan and Zeng (2023); Serapio-García et al. (2023); La Cava et al. (2024). Li et al. (2023) introduced Unsupervisedly-Built Personalized Lexicons (UBPL) to adjust Big Five traits during decoding, avoiding fine-tuning but risking training data bias. Similarly, Weng et al. (2024) and Dan et al. (2024) proposed ControlLM and P-tailor to efficiently simulate traits using control vectors, though these methods add complexity and may struggle with scalability. Finally, Mao et al. (2023) employed knowledge editing techniques like IKE, MEND, SERAC, and Prompt to manipulate traits like agreeableness, but MEND and SERAC still yield inconsistent results.
@@ 3 Methodology
This paper explores manipulation in autoregressive transformer models using Parameter-Efficient Fine-Tuning (PEFT) and In-Context Knowledge Editing (IKE), focusing on LLaMA-2-7B-Chat Touvron et al. (2023), LLaMA-3-8B-Instruct MetaAI (2023), and Mistral-7B-Instruct Jiang et al. (2023a).
3.1 Personality Dataset and Metric Classifier
@@
This work expands on the dataset from Mao et al. (2023), which consisted of opinion-based QA on specific topics, by better aligning with the topics and more accurately capturing personality traits. This study adds Openness and Conscientiousness to the original traits of Extraversion, Agreeableness, and Neuroticism. While Mao et al. (2023) excluded these dimensions, considering them similar to Agreeableness in generating viewpoints, we argue their inclusion is vital for a comprehensive analysis and understanding of trait influence on opinions. Adding Openness and Conscientiousness allows us to capture additional aspects of personality, such as intellectual curiosity, preference for novelty, and diligent behaviour, which are not fully encompassed by Agreeableness alone. The dataset contains $5000$ instances of GPT-3.5-based model generated opinion texts, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait).  [^1] Dataset: https://huggingface.co/datasets/holistic-ai/personality_manipulation.Instances were created using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression. The generated text was analysed using word clouds and text analysis to identify key linguistic patterns, thematic elements, and ensure lexical diversity associated with the Big Five personality traits, as detailed in Appendix A.1. Additionally, the text was manually verified to ensure alignment with the intended traits, providing a robust representation of trait-specific language usage.
We further developed a multi-class personality classifier using RoBERTa Liu et al. (2019), which was fine-tuned on personality dataset and achieved an accuracy of 91.9% on the test set.  [^2] Classifier: https://huggingface.co/holistic-ai/personality_classifier. Classifier validation involved human verification of textual feature importance using SHAP Lundberg and Lee (2017) and LIME Ribeiro et al. (2016) to ensure predictions aligned with human understanding. For instance, in the sentence, "I believe the First Indochina War had its consequences, paving the way for the withdrawal of French colonial forces. However, there were many factors at play, and it’s important to acknowledge the contributions of everyone involved." certain words play a pivotal role in predicting the Agreeableness trait as observed in Figures 51 and 52. Terms such as "contributions", "acknowledge", and "believe" have a strong positive contribution to the prediction of Agreeableness, as they suggest inclusiveness, recognition, and a conciliatory tone, which are characteristic of Agreeableness. On the other hand, words like "consequences" contribute negatively to the Agreeableness prediction because it often connotes conflict, repercussions, or negative outcomes, whereas Agreeableness is characterized by cooperation, empathy, and a focus on harmony and positive social interactions Liu and Sun (2020).
@@
Thus, as seen above and in other examples inA.12, the results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity. Through above analysis, we can conclude that the personality classifier has successfully captured the expected patterns from perspective of feature attribution. Further details about the classifier are in A.2.
3.2 Personality Manipulation Methods
The In-Context Knowledge Editing (IKE) method Zheng et al. (2023) was used as a baseline to manipulate embedded knowledge in LLMs, serving as a comparative foundation for evaluating prompt-based versus fine-tuning techniques in personality manipulation. The same prompt from Mao et al. (2023), provided in A.11, ensured consistency in comparing IKE with PEFT. We excluded methods like MEND due to inconsistent and gibberish outputs.
Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), was selected to reduce computational cost while maintaining performance. The process began by preparing the Personality dataset, formatting an additional column as <s>[INST] Question [/INST] Answer </s>. Models were loaded from Hugging Face using 4-bit quantization via BitsAndBytes Dettmers et al. (2023b) for efficient processing and storage. The temperature was set to 1. The hyperparameters, chosen for their balance of efficiency and performance Houlsby et al. (2019); Dettmers et al. (2024), are detailed in A.7. After training, fine-tuned LoRa weights were merged with the original model for evaluation.
3.3 Personality Manipulation Metrics
To evaluate the effectiveness of personality manipulation, two metrics were used: Trait Alignment (TA) and Personality Adjective Evaluation (PAE). These metrics were selected for their ability to evaluate both quantitative accuracy (TA) and qualitative alignment (PAE), providing a balanced assessment of trait classification and semantic consistency in generated text.
TA is computed using the metric classifier developed in this study by comparing predicted labels $\hat{y}_{i}$, which are obtained after classifying the generated responses, with true labels $y_{i}$, which correspond to the original target personality traits from the dataset. For a dataset with $N$ instances, TA is given by $\text{TA}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}(\hat{y}_{i}=y_{i})$, where $\mathbb{1}(\hat{y}_{i}=y_{i})$ is 1 if the prediction matches the true label and 0 otherwise, providing an average alignment between predictions and the intended personality traits.
PAE, inspired by Mao et al. (2023), uses Chain of Thought (CoT) prompting Wei et al. (2022), where a larger LLM (here GPT-4 Brown (2020)) scores generated text on a 1-5 scale based on its alignment with the target trait. The PAE score is calculated as the difference between the score of generated text and original text from the dataset, with higher difference indicating that the generated text after PEFT more effectively captures the intended personality traits compared to the original text. The final PAE is the mean of these score differences across all instances, $\text{PAE}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, where $s_{i}$ is the score difference for each instance $i$. Details of the prompt are provided in A.9.
3.4 Explainability Analysis
After PEFT, LLaMA-2-7B-Chat and Mistral-7B-Instruct models started to incorporate emojis in responses, despite no emojis being present in the PEFT data. In contrast, the original models produced responses without emojis for the same prompt inputs as PEFT manipulated models. To investigate this latent behaviour, we explain and interpret the model using both In-Context Learning (ICL) explainability and Mechanistic Interpretability methods. Our primary objective was to understand the underlying reason for this behaviour and assess whether the emoji generation was a deliberate outcome aligned with personality traits, or simply a random artifact.
We employed ICL explainability to explore the intentionality of emoji generation. Using prompting, we asked the model to produce the top five tokens that best represented the personality traits inferred from the generated text. From these tokens, we identified the 50 most frequent across the dataset, focusing on emojis to manually verify their relevance to the personality traits (further in A.3). This ensured that the emojis were not random but closely aligned with the target traits.
Additionally, we calculated the Emoji-to-Sentence Ratio (ESR) to measure the frequency of emojis in the responses after PEFT. The ESR was defined as: ESR=# Sentences with emojis# Total $\text{ESR}=\frac{\text{\# Sentences with emojis}}{\text{\# Total sentences}}.$ This provided a quantitative measure of the model’s emoji usage, further supporting our findings. We hypothesized that the emoji generation could stem from pre-training on diverse corpora containing emoji patterns Radford et al. (2019), with PEFT manipulation amplifying these emerging behaviours.
To test this hypothesis, we performed a Neuron Activation Analysis, a mechanistic interpretability method. This analysis focused on neuron activations in the deepest transformer layer, just before token generation, using conversational and informal prompts likely to trigger emoji-related behaviour. In this paper, we used ’Hey! It’s been a busy day for everyone. I hope you’re feeling good about everything [Uncaptioned image].’ as the conversational and informal prompt. This prompt was chosen for its Neutral and Open-Ended Tone. Moreover, The "[Uncaptioned image]" smile is friendly but subdued, allowing for a wide range of reactions without pushing for strong positivity or negativity Shi (2022). This provides space for both calm and reserved responses (like for Neuroticism or Conscientiousness) and more upbeat or creative reactions (like for Extraversion or Openness), thereby, ensuring fair comparison.
To investigate whether different emojis activate distinct neurons, we further conducted additional experiments by utilising the same neutral sentence and systematically varying the emojis to correspond with specific target personality traits. For example, we used [Uncaptioned image] for Agreeableness and [Uncaptioned image] for Neuroticism to observe whether the activations differed based on the emoji used. Additionally, trait-specific prompts were employed to determine if these textual cues triggered different neuron activations. In total, we used 17 prompts (as in A.6 and A.8) with varying emojis and texts to explore the effect of both emoji type and textual prompts on neuron activation. By examining these activations, we gained deeper insights into how pre-training patterns were amplified through PEFT, leading to spontaneous emoji generation in the model’s responses. To further support our claim that PEFT amplifies latent behaviours in LLMs, we calculated token probability of the "[Uncaptioned image]" emoji being generated as the next token using the same sentence above. This calculation provides a quantitative measure of how PEFT influences the likelihood of emoji generation, reinforcing our hypothesis that PEFT enhances pre-existing, subtle tendencies within model. The probabilities are provided in Appendix A.4.
4 Results and Discussion
4.1 Personality Manipulation
The performance of PEFT and IKE is presented in Figure 4 and detailed in Table 12 in A.10. Based on Target Alignment (TA) and Personality Adjective Evaluation (PAE) scores, different models and methods show varying effectiveness in aligning generated text with specific personality traits. In the LLaMA-2-7B-chat model, PEFT demonstrates strong trait alignment for Neuroticism and Extraversion but struggles with nuanced expressions, as indicated by negative PAE scores. Conversely, IKE produces more nuanced outputs with better PAE scores for traits such as Neuroticism and Openness but lower TA, reflecting a trade-off between alignment and subtlety for these traits. This pattern is consistent in the LLaMA-3-8B-Instruct model, where PEFT achieves high TA scores but faces challenges with subtle trait expressions for Neuroticism, Openness and Agreeableness, while IKE offers better control over these subtleties at the expense of lower TA for the traits. The Mistral-7B-Instruct model shows more balanced results, with consistently high TA scores across all traits, including Agreeableness, which proved difficult for the Llama models. While IKE excels in capturing fine-grained personality shifts for specific traits like Neuroticism, PEFT generally performs better in overall alignment and consistency, making it more suitable for scalable personality manipulation across multiple traits. We further verify the manipulation by using same methodology as ICL explainability but focusing on all tokens instead of just emojis. Through this approach (as in A.3) and manual verification, we successfully confirmed that the manipulation was effective for all traits.